iosangel:

Multiply i+√3 by its complex conjugate.

snowflake0531:

a+bi the conjugate would be a-bi so it would be \[(\sqrt{3}+i) \times (\sqrt{3}-i)\]

snowflake0531:

\[\sqrt{3}^2 -i\sqrt{3}+i\sqrt{3}-i^2 \\ 3 -(-1)\\3+1\\4\]